OpenStudy (rose21):

What value of x makes the equation true? two-eighths equals twelve over x 2 over 8 = 12 over x CHOICES: A:96 B:48 C:18 D:14

OpenStudy (anonymous):

\[\frac{ 2 }{ 8 }=\frac{ 12 }{ x }\]

OpenStudy (anonymous):

\[x=\frac{ 8*12 }{ 2 }=48\]

OpenStudy (rose21):

so first u have to multiply then divide?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you can divide 8 with 2 or 12 with 2.. but not both. and then multiply and you will have the same answer : 48

OpenStudy (anonymous):

\[\frac{ 8 }{ 2 }*12=48\]

OpenStudy (rose21):

ok thank you sooo much :D

OpenStudy (anonymous):

yw